# scVGAE: A Novel Approach using ZINB-Based Variational Graph Autoencoder for Single-Cell RNA-Seq Imputation

### v1

@@ 1 Introduction
    We developed a novel approach for scRNA-Seq imputation, utilizing a variational graph convolutional autoencoder with ZINB loss. This approach maintains cell-cell similarity while effectively fitting gene expression data into the distribution.
@@
    To enhance clustering stability, we utilize normalized matrices to maintain the original topology and prevent overfitting.
@@ 3.1 Data Preprocessing
$\begin{split}X&=\log{(X+1)}\\ X_{ij}&=\sqrt{\frac{X_{ij}}{\Sigma_{j}X_{ij}}}.\end{split}$ (1)
According to equation 1, the data is rescaled by log + 1 transformation, library normalization, and square root transformation.
@@ 3.4 ZINB-Based Graph Autoencoder
In addition, We integrated the normalized gene expression matrix into the imputation matrix $\hat{X}$ for each cell and gene. This integration helps retain the unique characteristics of genes and cells that might otherwise be lost during the imputation process. The formula is as follows:
$\hat{X}=\hat{X}+BatchNorm(X)+BatchNorm(X^{T})^{T}.$
By employing Batch Normalization, the data becomes trainable, which aids in maintaining specific representations and preventing overfitting.
Then $\hat{X}$ is utilized for the calculation of the reconstruction loss $L_{r}$ as follows:
@@ 4.4 Clustering performance
Metric
Dataset
scVGAE
Original
MAGIC
ALRA
DeepImpute
DCA
GNNImpute
ARI
Baron
0.7278
0.5606
0.7657
0.4235
0.8623
0.6392
0.5165
Brosens
0.5605
0.5083
0.4685
0.4967
0.1991
0.3775
0.3746
Carey
0.7978
0.6598
0.7437
0.6510
0.3750
0.4799
0.7023
cbmc
0.5937
0.4358
0.6622
0.4699
0.3348
0.4432
0.5632
Chang
0.1840
0.1747
0.1620
0.1351
0.1217
0.1502
0.0956
Fujii
0.4792
0.3565
0.4439
0.3462
0.1403
0.1978
0.3041
hcabm40k
0.0940
0.0371
0.0347
0.0895
0.0246
0.0352
0.0370
Hrvatin
0.7914
0.7119
0.7771
0.6021
0.4492
0.6681
0.6282
Jakel
0.6108
0.5043
0.4117
0.4419
0.3589
0.4634
0.4562
Jiang
0.2811
0.3421
0.2871
0.3435
0.1379
0.2612
0.2284
Manno
0.3095
0.2929
0.2123
0.2858
0.1269
0.1726
0.1883
Mingyao
0.2373
0.2303
0.1778
0.2020
0.1629
0.1855
0.1735
pbmc3k
0.6933
0.6381
0.6501
0.6370
0.2911
0.5122
0.5583
Xu
0.7357
0.6749
0.6995
0.6629
0.3445
0.4711
0.6531
AMI
Baron
0.7896
0.7500
0.8356
0.7143
0.7691
0.7567
0.6641
Brosens
0.6687
0.6355
0.6494
0.6296
0.3557
0.5506
0.5434
Carey
0.8146
0.7213
0.7782
0.7088
0.5577
0.6164
0.7266
cbmc
0.6855
0.6367
0.7462
0.6437
0.5244
0.6156
0.6258
Chang
0.3244
0.3135
0.3023
0.2558
0.2266
0.2855
0.1941
Fujii
0.5716
0.4767
0.5658
0.4749
0.2289
0.3216
0.4276
hcabm40k
0.1343
0.0525
0.0517
0.1311
0.0420
0.0532
0.0498
Hrvatin
0.8590
0.8514
0.8677
0.7979
0.6888
0.8155
0.6856
Jakel
0.6627
0.6316
0.5875
0.5986
0.4643
0.5718
0.5742
Jiang
0.4789
0.5101
0.4964
0.5360
0.2456
0.4017
0.3974
Manno
0.4735
0.4466
0.4301
0.4574
0.2995
0.3779
0.3680
Mingyao
0.3764
0.3224
0.2942
0.3105
0.2555
0.3039
0.2646
pbmc3k
0.7732
0.7248
0.7649
0.7389
0.4693
0.6661
0.6615
Xu
0.7929
0.7396
0.7814
0.7414
0.4942
0.5811
0.7011

### v2

@@ 1 Introduction
    Novel ZINB-Based Graph Autoencoder Approach: Integrates Zero-Inflated Negative Binomial (ZINB) loss with Graph Convolutional Networks (GCN) for scRNA-seq imputation, preserving cell-cell similarity while fitting gene expression data to appropriate distributions.
@@
    Compound Loss Function: Combines ZINB loss with reconstruction loss, maintaining original data distribution while effectively imputing missing values, offering a more flexible and efficient learning process compared to single-loss approaches.
  3. 3.
    Comprehensive Performance Evaluation: Demonstrates robustness and superiority of the proposed method through extensive testing on 14 diverse real-world scRNA-seq datasets, outperforming state-of-the-art methods in various cell clustering metrics.
@@ 3.1 Data Preprocessing
$\begin{split}X_{ij}&=\sqrt{\frac{X_{ij}}{\Sigma_{j}X_{ij}}}.\end{split}$ (1)
According to equation 1, the data is rescaled by library normalization, and square root transformation.
@@ 3.4 ZINB-Based Graph Autoencoder
In our previous work [13], we proposed a method to integrate normalized gene expression matrices into the imputation process. This approach helps preserve the unique characteristics of genes and cells during imputation:
$\hat{X}=\hat{X}+BatchNorm(X)+BatchNorm(X^{T})^{T}$ (5)
where $\hat{X}$ is the imputed matrix, $X$ is the original matrix, and $BatchNorm$ represents Batch Normalization. This normalization makes the data trainable, maintaining specific representations and preventing overfitting. The reconstruction loss $L_{r}$ is then calculated as:
@@ 4.4 Clustering performance
Metric Dataset scVGAE Original MAGIC ALRA DeepImpute DCA GNNImpute
ARI Baron 0.7278 0.5606 0.7657 0.4235 0.8623 0.6392 0.5165
Brosens 0.5605 0.5083 0.4685 0.4967 0.1991 0.3775 0.3746
Carey 0.7978 0.6598 0.7437 0.6510 0.3750 0.4799 0.7023
cbmc 0.5937 0.4358 0.6622 0.4699 0.3348 0.4432 0.5632
Chang 0.1840 0.1747 0.1620 0.1351 0.1217 0.1502 0.0956
Fujii 0.4792 0.3565 0.4439 0.3462 0.1403 0.1978 0.3041
hcabm40k 0.0940 0.0371 0.0347 0.0895 0.0246 0.0352 0.0370
Hrvatin 0.7914 0.7119 0.7771 0.6021 0.4492 0.6681 0.6282
Jakel 0.6108 0.5043 0.4117 0.4419 0.3589 0.4634 0.4562
Jiang 0.2811 0.3421 0.2871 0.3435 0.1379 0.2612 0.2284
Manno 0.3095 0.2929 0.2123 0.2858 0.1269 0.1726 0.1883
Mingyao 0.2373 0.2303 0.1778 0.2020 0.1629 0.1855 0.1735
pbmc3k 0.6933 0.6381 0.6501 0.6370 0.2911 0.5122 0.5583
Xu 0.7357 0.6749 0.6995 0.6629 0.3445 0.4711 0.6531
AMI Baron 0.7896 0.7500 0.8356 0.7143 0.7691 0.7567 0.6641
Brosens 0.6687 0.6355 0.6494 0.6296 0.3557 0.5506 0.5434
Carey 0.8146 0.7213 0.7782 0.7088 0.5577 0.6164 0.7266
cbmc 0.6855 0.6367 0.7462 0.6437 0.5244 0.6156 0.6258
Chang 0.3244 0.3135 0.3023 0.2558 0.2266 0.2855 0.1941
Fujii 0.5716 0.4767 0.5658 0.4749 0.2289 0.3216 0.4276
hcabm40k 0.1343 0.0525 0.0517 0.1311 0.0420 0.0532 0.0498
Hrvatin 0.8590 0.8514 0.8677 0.7979 0.6888 0.8155 0.6856
Jakel 0.6627 0.6316 0.5875 0.5986 0.4643 0.5718 0.5742
Jiang 0.4789 0.5101 0.4964 0.5360 0.2456 0.4017 0.3974
Manno 0.4735 0.4466 0.4301 0.4574 0.2995 0.3779 0.3680
Mingyao 0.3764 0.3224 0.2942 0.3105 0.2555 0.3039 0.2646
pbmc3k 0.7732 0.7248 0.7649 0.7389 0.4693 0.6661 0.6615
Xu 0.7929 0.7396 0.7814 0.7414 0.4942 0.5811 0.7011